# Sharp Thresholds for Factors in Random Graphs

### v1

@@ 5.2 Main Result
Let $\mathcal{B}_{1}$ be the event that the maximum degree of ${H_{F}}(n,\pi)$ is at least $\Delta=n^{\varepsilon}+\ln(n)n^{\varepsilon/2}$. Let $\mathcal{B}_{2}$ be the event that ${H_{F}}(n,\pi)$ contains an avoidable configuration with at most $2e(F)^{2}$ $F$-edges. Let $\mathcal{B}_{3}$ be the event that the coupling of $\mathcal{C}_{1}$ and $\mathcal{C}_{2}$ fails. Finally, let $\mathcal{B}=\mathcal{B}_{1}\cup\mathcal{B}_{2}\cup\mathcal{B}_{3}$. Note that $\mathcal{B}_{1}$ does not hold whp by the Chernoff bound, $\mathcal{B}_{2}$ does not hold whp by Lemma 3.2 (for small enough $\varepsilon>0$) and $\mathcal{B}_{3}$ does not hold whp by Proposition 4.4, thus $\mathcal{B}$ does not hold whp. Also, recall from the proof of Lemma 4.1 that overlapping clean $F$-cycles are avoidable configurations, so the clean $F$-cycles $\mathcal{C}_{1}=\mathcal{C}_{2}$ are pairwise vertex-disjoint unless $\mathcal{B}$ holds.
@@ 5.3 Conditional Probabilities
We follow [24] with the adjustments from [13]. Assume that the coupling fails for the first time in step $j$. Note that prior to $j$ we have chosen to include an $F$-edge $F_{i}$ in ${H_{F}}(n,\pi)$ exactly if we chose to include it in $G^{*}(n,p)$. Let $Y\subset[j-1]$ be the steps where we made these decisions. For all steps $i\in[j-1]\setminus Y$ we decided to exclude $F_{i}$ from ${H_{F}}(n,\pi)$, but only for a subset $N^{\prime}\subset[j-1]\setminus Y$ of steps $i\in N^{\prime}$ we decided to exclude $F_{i}$ from $G^{*}(n,p)$. For the remaining steps $[j-1]\setminus(Y\cup N^{\prime})$, we made no decision regarding the containment of $F_{i}$ in $G^{*}(n,p)$.
@@
Further, let $H_{0}$ be the $F$-graph on vertices $[n]$ with $F$-edges $F_{i}$ for $i\in Y$ and clean $F$-cycles $\mathcal{C}_{1}$. If $H_{0}\cup F_{j}$ has maximum degree at least $\Delta$, then $\mathcal{B}_{1}$ holds and we are done. Further, enumerate all clean $F$-cycles except those in $\mathcal{C}_{1}$ with $N^{\prime\prime}=\{F_{-1},\dots,F_{-\ell}\}$ and let $N=N^{\prime}\cup N^{\prime\prime}$.
@@
Before we turn to the core of the proof, we want to build some intuition. Note that $\pi^{\prime}_{i}=1$ exactly if $F_{i}$ is in some cycle in $\mathcal{C}_{1}$, in which case we also have $\pi_{i}=1$ and the coupling algorithm did choose to include $F_{i}$ almost surely in both models. Further, note that $\pi^{\prime}_{i}=0$ exactly if $F_{i}$ is in some cycle $C\in N^{\prime\prime}$ and all other $F$-edges $(F_{s})_{s\in S}$ of $C$ were previously included, i.e. $s\in Y\cap[i-1]$, or are covered by some cycle $C^{\prime}$ in $\mathcal{C}_{1}$, i.e. $F_{s}\in C^{\prime}$. If $C$ is a dense clean $F$-cycle, then we also have $\pi_{i}=0$ and $F_{i}$ is excluded in $G^{*}(n,p)$. In the new case that $C$ is a sparse clean $F$-cycle, we do have $\pi_{i}>0$ because both $F$-edges might still be included in $G^{*}(n,p)$ just as long as the corresponding dummy edge is not. However, in this case our coupling flips the coin with probability $\pi^{\prime}_{i}/\pi_{i}=0$, meaning that $F_{i}$ will be excluded in ${H_{F}}(n,\pi)$ and not tested in $G^{*}(n,p)$ almost surely.
@@
Now we turn to the bound for $\pi^{\prime}_{i}$. Let $H^{\prime}$ be the random $F$-graph on vertices $[n]$ where each $F$-edge of $H_{0}$ is included, and every other $F$-edge is included independently with probability $\pi$. Let $\mathcal{L}_{2}$ be the event that the cycles in $N^{\prime\prime}$ are not included, then we have
$\pi^{\prime}_{i}=\mathbb{P}(F_{i}\textrm{ in }H^{\prime}|\mathcal{L}_{2})\leq\mathbb{P}(F_{i}\textrm{ in }H^{\prime})=\pi$ (3)
@@
unless $\pi^{\prime}_{i}=1$ (cf. above), since $\mathcal{L}_{2}$ is a down-set in the product probability space corresponding to $H^{\prime}$, while the inclusion of $F_{i}$ is an up-set.
@@
Next, we turn to the bound for $\pi_{j}$, as detailed in [24] and partially inspired by the proof of Janson’s inequality and a variation suggested by Lutz Warnke [25]. The following arguments are based on $G^{*}(n,p)$, so the discussion always involves dummy edges as well unless explicitly stated otherwise. Let $E_{i}$ be the edge set of $F_{i}$ for all $i$ (including any dummy edges) and let $R$ be the edge set of the union graph of $H_{0}$.
@@
Now, we split $Q=Q_{\mathrm{cb}}+Q_{\mathrm{cg}}+Q_{\mathrm{eb}}+Q_{\mathrm{eg}}$ for the discussion in the subsequent four sections, where $Q_{\mathrm{cb}}$ is given by the contributions of cycles, i.e. negative $i$, with $E_{i}\subset E_{j}\cup R$, $Q_{\mathrm{cg}}$ by the remaining contributions of cycles, $Q_{\mathrm{eb}}$ by the contributions of $F$-edges, i.e. positive $i$, with $E_{i}\subset E_{j}\cup R$, and $Q_{\mathrm{eg}}$ by the remaining contributions of $F$-edges.
@@ 5.4 Bad Edges
In this section, we argue that $Q_{\mathrm{eb}}=0$ unless $\mathcal{B}$ holds, in which case we are done. For $i\in N_{1}$ contributing to $Q_{\mathrm{eb}}$, we have $E_{i}\subset E_{j}\cup R$. With $i\in N_{1}$, we know that $F_{i}$ is not in $H_{0}\cup F_{j}$ and thus also not in ${H_{F}}(n,\pi)$. On the other hand, $H_{0}\cup F_{j}$ is contained in ${H_{F}}(n,\pi)$ and hence $E_{i}\subset E_{j}\cup R$ implies that $F_{i}$ is in the $F$-graph of the shadow graph of ${H_{F}}(n,\pi)$, meaning that $F_{i}$ is induced. By Lemma 3.1, there exists an avoidable configuration or a clean $F$-cycle inducing $F_{i}$. If the former is true, then $\mathcal{B}_{2}$ holds and we are done. Otherwise, it needs to be a cycle in $\mathcal{C}_{1}=\mathcal{C}_{2}$. But this implies that $\pi_{i}=1$, which in turn implies that $i\not\in N_{1}$, a contradiction.
@@ 5.5 Bad Cycles
In this section, we argue that $Q_{\mathrm{cb}}=0$ unless $\mathcal{B}$ holds, in which case we are done. For $i\in N_{1}$ contributing to $Q_{\mathrm{cb}}$, we have $E_{i}\subset E_{j}\cup R$. With $i\in N_{1}$, we know that $F_{i}$ is not in $H_{0}\cup F_{j}$ and thus also not in ${H_{F}}(n,\pi)$. On the other hand, $H_{0}\cup F_{j}$ is contained in ${H_{F}}(n,\pi)$ and hence $E_{i}\subset E_{j}\cup R$ suggests that $F_{i}$ is in the $F$-graph of the shadow graph of ${H_{F}}(n,\pi)$, meaning that some $F$-edges of $F_{i}$ have to be induced. As in Section 5.4 and using Lemma 3.1, if avoidable configurations are involved, then $\mathcal{B}_{2}$ holds and we are done.
@@ 5.6 Good Edges
For $i\in N_{1}$ contributing to $Q_{\mathrm{eg}}$, we have $E_{i}\setminus(E_{j}\cup R)\neq\emptyset$. On the other hand, by the definition of $N_{1}$, we have $(E_{i}\cap E_{j})\setminus R\neq\emptyset$. To obtain a bound on the number of vertex sets, we consider the graph intersection $S$ given by the vertex set of $F_{i}$ and the edges from $E_{j}\cup R$ that are in $F_{i}$. Hence, the contribution can be written as $p^{|E_{i}\setminus(E_{j}\cup R)|}=p^{e(F)-e(S)}$. Now, we count the number of contributions given $e(S)$ and $c(S)$. Knowing the vertex set, the factor $r!/\operatorname{aut}(F)$ bounds the number of contributions by $F_{i}$ on this vertex set. For the vertex set, we choose an anchor for each component, respecting that one component has to take one of the vertices in $E_{j}$, which gives $rn^{c(S)-1}$. Then, for each component individually, given the chosen anchor vertex, the remaining vertices of the component can be picked step-by-step from the neighbours (in $E_{j}\cup R$) of the vertices picked so far, yielding the bound $r(r-1)\Delta$ for each such step. Thus, the total number of vertex sets is bounded by $O(n^{c(S)-1}\Delta^{r-c(S)})=O(n^{c(S)-1+(r-c(S))\varepsilon})$. Thus, the total contribution for given $e(S)$ and $c(S)$ to $Q_{\mathrm{eg}}$ is bounded by $O(n^{f(S)})$, where
@@
Note that $f_{1}$ is additive over disjoint graphs, zero for isolated vertices and for each non-trivial component $S^{\prime}$ we have $d_{1}(S^{\prime})<d_{1}(F)$. This shows that $f(S)<0$ for any of these subgraphs $S$ of $F$ if $\varepsilon$ is small enough, yielding that the contribution from $Q_{\mathrm{eg}}$ is $o(1)$.
@@ 5.7 Good Cycles
For $i\in N_{1}$ contributing to $Q_{\mathrm{cg}}$ where $F_{i}$ is a copy of $C$, we have $E_{i}\setminus(E_{j}\cup R)\neq\emptyset$. On the other hand, by the definition of $N_{1}$, we have $(E_{i}\cap E_{j})\setminus R\neq\emptyset$. Let $S$ be given by the vertex set of $F_{i}$ and the edges from $E_{j}\cup R$ on these vertices, so the contribution can be written as $p^{|E_{i}\setminus(E_{j}\cup R)|}=p^{e(G)-e(S)}$, where $G$ is the clean d-cycle corresponding to $C$. Now, we count the number of vertex sets given $e(S)$, $c(S)$ and the type $C$. As before, we choose an anchor for each component, respecting that one component has to take one of the vertices in $E_{j}$, which gives $O(n^{c(S)-1})$. Then, for each component individually, given the anchor vertex we may choose the remaining vertices step-by-step from the neighbours of the vertices we have chosen so far. Thus, the total is bounded by $O(n^{c(S)-1}\Delta^{v(S)-c(S)})=O(n^{c(S)-1+(v(S)-c(S))\varepsilon})$. Thus, the total contribution for given $e(S)$ and $c(S)$ to $Q_{\mathrm{eg}}$ is bounded by $O(n^{g(S)})$, where
@@
Now, we show that $g_{1}(S)<0$ for all $S$. Note that $g_{1}(S)=f_{1}(S)-1$ with $f_{1}$ from Section 5.6. Since $f_{1}$ is additive over disjoint graphs and, as in Equation (2), $d_{1}(S^{\prime})\leq d_{1}(F)$ for any component $S^{\prime}$ of $S$ which does not cover all overlap vertices of $F_{i}$, meaning $f_{1}(S^{\prime})\leq 0$, we may restrict to the components which do cover all overlaps. But clearly, there can be only one such component, so we may assume that $S$ is connected and covers all overlaps. But then we have
$g_{1}(S)=\left(\frac{e(S)}{v(S)d_{1}(F)}-1\right)v(S)$
and hence Lemma 3.3 completes the proof that $g_{1}(S)<0$.
@@
As indicated throughout the proof, we now choose any
$\max\left(\max_{S}f_{1}(S),\max_{S}g_{1}(S)\right)<\delta<0$
and $\varepsilon$ sufficiently small and in particular such that $f(S),g(S)<\delta$ over all $S$. Then, after collecting the contributions from all four cases, we have

### v2

@@ 5.2 Main Result
Let $\mathcal{B}_{1}$ be the event that the maximum degree of ${H_{F}}(n,\pi)$ is at least $\Delta=n^{\varepsilon}+\ln(n)n^{\varepsilon/2}$. Let $\mathcal{B}_{2}$ be the event that ${H_{F}}(n,\pi)$ contains an avoidable configuration with at most $2e(F)^{2}$ $F$-edges. Let $\mathcal{B}_{3}$ be the event that the coupling of $\mathcal{C}_{1}$ and $\mathcal{C}_{2}$ fails. Finally, let $\mathcal{B}=\mathcal{B}_{1}\cup\mathcal{B}_{2}\cup\mathcal{B}_{3}$. Note that ${\mathbb{P}}(\mathcal{B}_{1})=o(1)$ by the Chernoff bound, ${\mathbb{P}}(\mathcal{B}_{2})=o(1)$ by Lemma 3.2 (for small enough $\varepsilon>0$) and ${\mathbb{P}}(\mathcal{B}_{3})=o(1)$ by Proposition 4.4, thus whp $\mathcal{B}$ does not hold. Also, recall from the proof of Lemma 4.1 that overlapping clean $F$-cycles are avoidable configurations, so the clean $F$-cycles $\mathcal{C}_{1}=\mathcal{C}_{2}$ are pairwise vertex-disjoint unless $\mathcal{B}$ holds.
@@ 5.3 Conditional Probabilities
We follow [24] with the adjustments from [13]. Assume that the coupling fails for the first time in step $j\geq 1$. Note that prior to $j$ we have chosen to include an $F$-edge $F_{i}$ in ${H_{F}}(n,\pi)$ exactly if we chose to include it in $G^{*}(n,p)$. Let $Y\subset[j-1]$ be the steps where we made these decisions. For all steps $i\in[j-1]\setminus Y$ we decided to exclude $F_{i}$ from ${H_{F}}(n,\pi)$, but only for a subset $N^{\prime}\subset[j-1]\setminus Y$ of steps $i\in N^{\prime}$ we decided to exclude $F_{i}$ from $G^{*}(n,p)$. For the remaining steps $j-1]\setminus(Y\cup N^{\prime})$, we made no decision regarding the containment of $F_{i}$ in $G^{*}(n,p)$.
@@
Further, let $H_{0}$ be the $F$-graph on vertices $[n]$ with $F$-edges $F_{i}$ for $i\in Y$ and clean $F$-cycles $\mathcal{C}_{1}$. If $H_{0}\cup F_{j}$ has maximum degree at least $\Delta$, then $\mathcal{B}_{1}$ holds and we are done. Further, enumerate all (potential) clean $F$-cycles except those in $\mathcal{C}_{1}$ with $\mathcal{C}_{1}^{c}=\{F_{-1},\dots,F_{-\ell}\}$, set $N^{\prime\prime}=\{-1,\dots,-\ell\}$ and let $N=N^{\prime}\cup N^{\prime\prime}$.
@@
Before we turn to the core of the proof, we want to build some intuition. Note that $\pi^{\prime}_{j}=1$ exactly if $F_{j}$ is in some cycle in $\mathcal{C}_{1}$, in which case we also have $\pi_{j}=1$ and the coupling algorithm did choose to include $F_{j}$ almost surely in both models. Further, note that $\pi^{\prime}_{j}=0$ exactly if $F_{j}$ is in some cycle $C\in\mathcal{C}_{1}^{c}$ and all other $F$-edges $(F_{s})_{s\in S}$ of $C$ were previously included, i.e. $s\in Y\cap[j-1]$, or are included in some cycle $C^{\prime}$ in $\mathcal{C}_{1}$, i.e. $F_{s}\in C^{\prime}$. If $C$ is a dense clean $F$-cycle, then we also have $\pi_{j}=0$ and $F_{j}$ is excluded in $G^{*}(n,p)$. In the new case that $C$ is a sparse clean $F$-cycle, we do have $\pi_{j}>0$ because both $F$-edges might still be included in $G^{*}(n,p)$ just as long as the corresponding dummy edge is not. However, in this case our coupling flips the coin with probability $\pi^{\prime}_{j}/\pi_{j}=0$, meaning that $F_{j}$ will be excluded in ${H_{F}}(n,\pi)$ and not tested in $G^{*}(n,p)$ almost surely.
@@
Now we turn to the bound for $\pi^{\prime}_{j}$. Let $H^{\prime}$ be the random $F$-graph on vertices $[n]$ where each $F$-edge of $H_{0}$ is included, and every other $F$-edge is included independently with probability $\pi$. Let $\mathcal{L}_{2}$ be the event that the cycles in $\mathcal{C}_{1}^{c}$ are not included, then we have
$\pi^{\prime}_{j}=\mathbb{P}(F_{j}\textrm{ in }H^{\prime}|\mathcal{L}_{2})\leq\mathbb{P}(F_{j}\textrm{ in }H^{\prime})=\pi$ (5)
@@
unless $\pi^{\prime}_{j}=1$ (cf. above), since $\mathcal{L}_{2}$ is a down-set in the product probability space corresponding to $H^{\prime}$, while the event $\{F_{j}\in H^{\prime}\}$ is an up-set.
@@
Next, we turn to the bound for $\pi_{j}$, as detailed in [24] and partially inspired by the proof of Janson’s inequality and a variation suggested by Lutz Warnke [25]. The following arguments are based on $G^{*}(n,p)$, so the discussion always involves dummy edges as well unless explicitly stated otherwise. Let $E_{i}$ be the edge set (including any dummy edges) of $F_{i}$ for all $i$ (positive and negative) and let $R$ be the edge set of the shadow graph of $H_{0}$.
@@
Now, we split $Q=Q_{\mathrm{cb}}+Q_{\mathrm{cg}}+Q_{\mathrm{eb}}+Q_{\mathrm{eg}}$ for the discussion in the subsequent four sections, as follows:
  • $Q_{\mathrm{cb}}$ is the contribution coming from bad cycles, i.e. $i<0$ with $E_{i}\subset E_{j}\cup R$
  • $Q_{\mathrm{cg}}$ is the contribution coming from good cycles, i.e. $i<0$ with $E_{i}\not\subset E_{j}\cup R$
  • $Q_{\mathrm{eb}}$ is the contribution coming from bad edges, i.e. $i>0$ with $E_{i}\subset E_{j}\cup R$
  • $Q_{\mathrm{eg}}$ is the contribution coming from good edges, i.e. $i>0$ with $E_{i}\not\subset E_{j}\cup R$.
@@ 5.4 Bad Edges
In this section, we argue that $Q_{\mathrm{eb}}=0$ unless $\mathcal{B}$ holds, in which case we are done. For $i\in N_{1}$ contributing to $Q_{\mathrm{eb}}$, we have $E_{i}\subset E_{j}\cup R$. With $i\in N_{1}$, we know that $F_{i}$ is not in $H_{0}\cup F_{j}$ and thus also not in ${H_{F}}(n,\pi)$. On the other hand, $H_{0}\cup F_{j}$ is contained in ${H_{F}}(n,\pi)$ and hence $E_{i}\subset E_{j}\cup R$ implies that $F_{i}$ is in the $F$-graph of the shadow graph of ${H_{F}}(n,\pi)$, meaning that $F_{i}$ is induced. By Lemma 3.1, there exists an avoidable configuration or a clean $F$-cycle inducing $F_{i}$. If the former is true, then $\mathcal{B}_{2}$ holds and we are done. Otherwise, $F_{i}$ needs to be induced by a cycle in $\mathcal{C}_{1}=\mathcal{C}_{2}$. But this implies that $\pi_{i}=1$, which in turn implies that $i\not\in N_{1}$, a contradiction.
@@ 5.5 Bad Cycles
In this section, we argue that $Q_{\mathrm{cb}}=0$ unless $\mathcal{B}$ holds, in which case we are done. For $i\in N_{1}$ contributing to $Q_{\mathrm{cb}}$ (recall that this means that $i<0$ and $F_{i}$ is a clean cycle), we have $E_{i}\subset E_{j}\cup R$. With $i\in N_{1}$, we know that $F_{i}$ is not in $H_{0}\cup F_{j}$ and thus also not in ${H_{F}}(n,\pi)$. On the other hand, $H_{0}\cup F_{j}$ is contained in ${H_{F}}(n,\pi)$ and hence $E_{i}\subset E_{j}\cup R$ suggests that $F_{i}$ is in the $F$-graph of the shadow graph of ${H_{F}}(n,\pi)$, meaning that some $F$-edges of $F_{i}$ have to be induced. As in Section 5.4 and using Lemma 3.1, if avoidable configurations are involved, then $\mathcal{B}_{2}$ holds and we are done.
@@ 5.6 Good Edges
For $i\in N_{1}$ contributing to $Q_{\mathrm{eg}}$, we have $E_{i}\setminus(E_{j}\cup R)\neq\emptyset$. On the other hand, by the definition of $N_{1}$, we have $(E_{i}\cap E_{j})\setminus R\neq\emptyset$. To obtain a bound on the number of vertex sets, we consider the graph intersection $S$ given by the vertex set of $F_{i}$ and the edges from $E_{j}\cup R$ that are in $F_{i}$, i.e. $S=(V(F_{i}),E_{i}\cap(E_{j}\cup R))$. Hence, the contribution can be written as $p^{|E_{i}\setminus(E_{j}\cup R)|}=p^{e(F)-e(S)}$. Now, we count the number of contributions given $e(S)$ and $c(S)$. We first count the number of vertex sets that lead to such $S$; to each vertex sets there are then at most $r!/\operatorname{aut}(F)$-many possible $F_{i}$. For the vertex set, we choose an anchor for each component, respecting that at least one component has to take one of the vertices in $E_{j}$, which gives $rn^{c(S)-1}$. Then, for each component individually, given the chosen anchor vertex, the remaining vertices of the component can be picked step-by-step from the neighbours (in $E_{j}\cup R$) of the vertices picked so far, yielding the bound $r(r-1)\Delta$ for each such step. Thus, the total number of vertex sets is bounded by $\mathcal{O}(n^{c(S)-1}\Delta^{r-c(S)})=\mathcal{O}(n^{c(S)-1+(r-c(S))\varepsilon})$. Thus, the total contribution for given $e(S)$ and $c(S)$ to $Q_{\mathrm{eg}}$ is bounded by $\mathcal{O}(n^{f(S)})$, where
@@
Note that $f_{1}$ is additive over disjoint graphs and zero for isolated vertices. For each non-trivial component $S^{\prime}$ of $S$ we have $d_{1}(S^{\prime})<d_{1}(F)$, implying $f_{1}(S^{\prime})=\left(\frac{d_{1}(S^{\prime})}{d_{1}(F)}-1\right)(v(S^{\prime})-1)<0$. This shows that $f(S)<0$ for any of these subgraphs $S$ of $F$ if $\varepsilon$ is small enough, yielding that the contribution from $Q_{\mathrm{eg}}$ is $o(1)$.
@@ 5.7 Good Cycles
For $i\in N_{1}$ contributing to $Q_{\mathrm{cg}}$ where $F_{i}$ is a copy of $C$, we have $E_{i}\setminus(E_{j}\cup R)\neq\emptyset$. On the other hand, by the definition of $N_{1}$, we have $(E_{i}\cap E_{j})\setminus R\neq\emptyset$. Let $S$ be given by the vertex set of $F_{i}$ and the edges from $E_{j}\cup R$ on these vertices, so the contribution can be written as $p^{|E_{i}\setminus(E_{j}\cup R)|}=p^{e(G)-e(S)}$, where $G$ is the clean d-cycle corresponding to $C$. Now, we count the number of vertex sets given $e(S)$, $c(S)$ and the type $C$. As in Section 5.6, we choose an anchor for each component, respecting that one component has to take one of the vertices in $E_{j}$, which gives $\mathcal{O}(n^{c(S)-1})$. Then, for each component individually, given the anchor vertex we may choose the remaining vertices step-by-step from the neighbours of the vertices we have chosen so far. Thus, the total is bounded by $\mathcal{O}(n^{c(S)-1}\Delta^{v(S)-c(S)})=\mathcal{O}(n^{c(S)-1+(v(S)-c(S))\varepsilon})$. Thus, the total contribution for given $e(S)$ and $c(S)$ to $Q_{\mathrm{eg}}$ is bounded by $O(n^{g(S)})$, where
@@
Now, we show that $g_{1}(S)<0$ for all $S$. Note that $g_{1}(S)=f_{1}(S)-1$ with $f_{1}$ from Section 5.6 and recall that $f_{1}$ is additive over disjoint graphs. Furthermore, as in Equation (3), we have $d_{1}(S^{\prime})\leq d_{1}(F)$ for any component $S^{\prime}$ of $S$ which does not cover all overlap vertices of the clean cycle $F_{i}$, which implies $f_{1}(S^{\prime})\leq 0$. It is thus sufficient to show $g_{1}(S^{\prime})<0$ for a component covering all overlap vertices (and such $S^{\prime}$ is unique in $S$, if it exists). For such $S^{\prime}$ we have
$g_{1}(S)=\left(\frac{e(S)}{v(S)d_{1}(F)}-1\right)v(S)<\left(\frac{d(F)}{d_{1}(F)}-1\right)v(S_{1})<0$
by Lemma 3.4, completing the proof that $g_{1}(S)<0$.
@@
As indicated throughout Sections 5.6 and 5.7, we now choose any $\delta>0$ such that
$\max\left(\max_{S}f_{1}(S),\max_{S}g_{1}(S)\right)<-2\delta<0$
and $\varepsilon>0$ sufficiently small as required throughout the proof, e.g. by ${\mathbb{P}}(\mathcal{B}_{2})=o(1)$, the proof of Proposition 4.4, and in particular such that $f(S),g(S)<-\delta$ over all $S$. Then, after collecting the contributions from all four cases, we have